Find the smallest positive integer N that satisfies all of the following conditions:


• N is a square.


• N is a cube.


• N is an odd number.


• N is divisible by twelve prime numbers.


How many digits does this number N have?


1
Expert's answer
2021-11-21T15:56:04-0500

Let;

"x² = N;\\
y³ = N\\
N = \textsf{odd}\\
\dfrac{N}{2×3×5×7×11×13×17×19×23×29×31×37}= z \ (\textsf{where z is an integer})\\
N=7420738134810z"


"\therefore" N is a multiple of 7420738134810


for N to be a square and a cube, and to be a multiple of a multiple of prime numbers, the multiple of prime numbers have to be exponentiated to (2×3)

"\therefore N = 7420738134810^6 = 166986990622241277447975946146208311466303689437571218130959563879576881000000"